This problem is taken from Ahlborn's Zoological Ph
Answer #1

(a) Using conservation of energy

(Potential energy at height) + (kinetic energy at height) = (Potential energy at ground) + (kinetic energy at ground)

(mgh) + (1/2m*02) = (mg*0) + (1/2 mUo2)

Uo = (2gh)1/2

Uo = (2 * 9.8 * 2)1/2

Uo = 6.26 m/s

-----------------------------------------------------------------------------------

(b) Using relation v2 - Uo2 = 2aΔy

0 - (6.26)2 = 2 a (0.05)

a = - 391.88 m/s2

----------------------------------------------------------------